OpenStudy (babalooo):

Fun question. 6÷2(1+2) is the answer 1 or 9? :D I know the answer... just for fun because I'm bored looking over other ppls's problems LOL

OpenStudy (anonymous):

9

OpenStudy (gschibby):

Nope, 1 ;)

OpenStudy (anonymous):

oh......

OpenStudy (anonymous):

9

OpenStudy (gschibby):

\[\frac{ 6 }{ 2(1+2)}=\frac{ 6 }{ 2*1+2*2 }=\frac{ 6 }{ 6 }=1\]

OpenStudy (anonymous):

oh......ok

OpenStudy (babalooo):

Actually the answer is 9 lol :D 6÷2(1+2) is not 6÷(2(1+2)) so we solve the brackets first like usual and we get 6÷2(3) which is also the same as 6÷2x3. We solve it from left to right so 3x3 = 9 ;D

OpenStudy (anonymous):

yeah!!!!

OpenStudy (babalooo):

\[6\div2(1+2) = 6\div2(3) = 6\div2\times3 = 9\]

OpenStudy (babalooo):

sorry i had to copy and paste the divide sign somewhere cuz they don't have all the pretty tools when having to to ask the question